OpenStudy (anonymous):

for f(x)=x^2-3x find the following f(b-3)

OpenStudy (anonymous):

explain how you got the answer also please

OpenStudy (anonymous):

just put in b-3 in place of x (b-3)^2 - 3(b-3) = b^2 - 6b + 9 - 3b -9 = b^2 -9b